# Some addition theorems for spin-weighted spherical harmonics

### v1

@@ I Introduction
The outline of this work is as follows. Sec. II reviews the key definitions and properties of SWSH which are required for our analysis. For ease of reference, our new addition theorems are presented in Sec. III, prior to their derivation in Sec. IV. Our conclusions can be found in Sec. V.
II Properties of SWSH
@@
where * denotes complex conjugation.
@@
Our results in this paper are valid for both integer and half-integer spins $s$. When $s$ is a half-integer, the factor of $(-1)^{s}$ in (1) will be purely imaginary (see, for example, [24] for some specific examples of SWSH in this case).
@@
The definitions (5) are, strictly speaking, valid only when $\phi-\phi^{\prime}$ is not a multiple of $\pi$. The corresponding definitions of the Euler angles when $\phi-\phi^{\prime}$ is a multiple of $\pi$ can be found in [15]. For our purposes, we require only that $\alpha=\beta=\gamma=0$ in the coincidence limit $\theta^{\prime}=\theta$, $\phi^{\prime}=\phi$.
@@
$\sum_{m=-\ell}^{\ell}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)=\frac{2\ell+1}{4\pi}\delta_{s,s^{\prime}},$ (9)
@@
III Further addition theorems
@@
Taking the coincidence limit $\theta^{\prime}=\theta$, $\phi^{\prime}=\phi$ in each of the results in (10) gives respectively, using (7):
@@
When $s=0$, the results (13) reduce to those found in, for example, App. C of [28] (see also [27, §5.10]).
@@
The precise form of the addition theorems (10, 12) depends on the conventions used for the definition of the SWSH . However, the final results in (13) are independent of the choice of phase in the SWSH .
IV Derivation of new addition theorems

### v2

@@ I Introduction
The outline of this work is as follows. In Sec. II we discuss some of the applications of SWSH in physics and motivate the addition theorems which are the focus of our work. Sec. III reviews the key definitions and properties of SWSH which are required for our analysis. For ease of reference, our new addition theorems are presented in Sec. IV, prior to their derivation in Sec. V. Our conclusions can be found in Sec. VI.
II Physical applications of SWSH addition theorems
SWSH typically arise in physical applications when one is considering a quantity (such as a field perturbation) which depends on spherical polar coordinates. Let us give a few examples:
Data analysis
Data from surveys such as those of the cosmic microwave background [6] or satellite gravity gradiometry [10] is naturally defined on a two-sphere. An expansion of observables in terms of SWSH facilitates computationally-efficient methods of analyzing such data [8, 7].
Modelling anisotropies
When physical systems are not spherically symmetric, an expansion in SWSH provides a powerful framework for the study of anisotropies, for example in fluid turbulence [17] or the cosmic microwave background [5], since the coefficients of the SWSH in such an expansion correspond to excitations of a particular multipole [18, 11].
Perturbations of black holes
Classical field equations on black hole backgrounds play an important role many aspects of general relativity, including the modelling of gravitational waves. If the background black hole is spherically symmetric, then separable mode solutions of the classical field equation involve SWSH (see, for example, [14]).
Our focus in this paper is addition theorems for SWSH . The theorems of the type we consider involve a product of two SWSH at different points over the two-sphere, summed over one of the quantum numbers of the SWSH (see Sec. III for more precise details). Such products of two SWSH at different points arise, for example, when one is interested in the Green function for a classical field equation on a flat or curved space-time. A Green function is a distributional solution of a classical field equation which depends on two points in space-time. In applications, it is often useful to have the Green function written as a sum over separable mode solutions of the field equation, see, for example, [11]. When the mode solutions involve SWSH , the Green function therefore involves a product of SWSH at different points on the two-sphere, and the quantum numbers in the SWSH label the field modes and are summed over. Addition theorems of the type considered in this paper are particularly useful in situations where the background on which the field propagates (such as a black hole space-time) is spherically symmetric, in which case the functions other than the SWSH in the field modes do not depend on one of the quantum numbers (in particular, the azimuthal quantum number $m$ as defined in Sec. III) in the SWSH . In this case, an addition theorem involving a sum over $m$ enables the Green function to be simplified.
We close this section by giving more details of an application of our addition theorems for SWSH in quantum field theory in curved space-time [27, 28, 29], in particular the computation of expectation values of observables such as the stress-energy tensor. The stress-energy tensor expectation value is of central importance in quantum field theory on curved space-time, since it governs, via the semiclassical Einstein equations, the backreaction of the quantum field on the curved space-time. To find the stress-energy tensor expectation value, one starts with the Green function for the classical equation of the particular field under consideration (which, as explained above, involves products of SWSH at different points on the two-sphere). One then applies a second-order differential operator to the Green function, with the result that one or more derivatives act on the SWSH in the sum. If one considers a field of nonzero spin on a four-dimensional black hole (see, for example, [28]), the resulting expectation values involve sums over SWSH of precisely the form we prove here (see Corollary 11 in Sec. IV below) for a particular value of the spin-weight, while a scalar field on a particular five-dimensional black hole [29] has a tower of modes involving SWSH of arbitrary (integer or half-integer) spin, with the result that we require the addition theorems in Corollary 11 for general spin-weight.
Having motivated the SWSH addition theorems which we will derive in this paper, in the next section we review the salient properties of SWSH , before turning to the addition theorems themselves.
III Properties of SWSH
@@
where * denotes complex conjugation. The Wigner $D$-matrices take the form [15]
$D^{\ell*}_{m,-s}(\phi,\theta,\chi)={\rm{e}}^{{\rm{i}}m\phi}d_{m,-s}^{\ell}(\theta){\rm{e}}^{-{\rm{i}}s\chi}$ (2)
where $d_{m,-s}^{\ell}(\theta)$ are real functions given by [15]
$d_{m,-s}^{\ell}(\theta)=\frac{(-1)^{\ell+m}}{2^{\ell}}{\sqrt{\frac{\left(\ell-m\right)!}{\left(\ell+m\right)!\left(\ell+s\right)!\left(\ell-s\right)!}}}\\ \times\left(1-\cos\theta\right)^{\frac{1}{2}\left(m+s\right)}\left(1+\cos\theta\right)^{\frac{1}{2}\left(m-s\right)}\\ \times\left(\frac{d}{d\left[\cos\theta\right]}\right)^{\ell+m}\left[\left(1-\cos\theta\right)^{\ell-s}\left(1+\cos\theta\right)^{\ell+s}\right].$ (3)
@@
Our results in this paper are valid for both integer and half-integer spins $s$. When $s$ is a half-integer, the factor of $(-1)^{s}$ in (1) will be purely imaginary (see, for example, [24] for some specific examples of SWSH in this case). The quantities ±mplus-or-$\ell\pm m$, ±$\ell\pm s$ and $m\pm s$ appearing in the Wigner $D$-matrices (3) are always integers.
@@
The definitions (7) are, strictly speaking, valid only when $\phi-\phi^{\prime}$ is not a multiple of $\pi$. In the coincidence limit $\theta^{\prime}\rightarrow\theta$, $\phi^{\prime}\rightarrow\phi$, taking the appropriate limit of (7) gives $\alpha=\beta=\gamma=0$.
@@
$\sum_{m=-\ell}^{\ell}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)=\frac{2\ell+1}{4\pi}\delta_{s,s^{\prime}}.$ (11)
@@
IV Further addition theorems
@@
Taking the coincidence limit $\theta^{\prime}=\theta$, $\phi^{\prime}=\phi$ in each of the results in (12) gives, respectively, using (9):
@@
When $s=0$, the results (15) reduce to those found in, for example, App. C of [27] (see also [30, §5.10]).
@@
The precise form of the addition theorems (12, 14) depends on the conventions used for the definition of the SWSH . However, the final results in (15) are independent of the choice of phase in the SWSH .
V Derivation of new addition theorems